Hi I need helping trying to determine if an equation is a function... I just don't understand
here are some examples:
x + 2y =4
x^2 + 5y^2 =1
x^2 + y-7 =0

Answer by stanbon(75887)

Solve each problem for y:
x + 2y =4
y = (-1/2)x+2
For each value of x there will be one value for y so it is a function.
----------------
x^2 + 5y^2 =1
5y^2=-x^2+1
y^2 = (1-x^2)/5
y= + sqrt[(1-x^2)/5] or y=-[sqrt[(1-x^2)/5]
For each value of x there will be two values for y so it is NOT a function.
----------------------
x^2 + y-7 =0
y = -x^2-7
For each value of x there will be one value for y so it is a function.
===============
Cheers,
Stan H.